Dewey is considering starting his own business as a photographer. Dewey is currently a tax attorney, making $150,000 per year and working 60 hours per week. Dewey estimates that it will take 6 hours to service each studio client and 18 hours for each bridal client He is limiting his work hours to the 60 hours he previously worked each week. He charges $500 per studio client and $5000 per bridal client. He spent $5,000 of his savings that had been earning 4% interest in order to purchase his camera and lighting equipment. He borrowed $10,000 from the bank to convert two rooms in his home to an office and studio space. He would also like you to determine what his annual revenue will be in addition to calculating the following:

1. What is Dewey's accounting profit?
2. Suppose Dewey could have used his talents to run a similar kind of business instead. If he values his entrepreneurial skill at $10,000 annually, find Dewey's total implicit costs.
3. What was Dewey's economic profit last year?
1
Expert's answer
2020-03-02T08:25:47-0500

1.Annual revenue= (5 000×3+500)×52=$ 806 000


2.accounting profit = 806 000 - 5 000 - 10 000 =$ 791 000


3.total implicit costs = |10 000- 150 000 - 5 000×0.04|= $ 140 200


4. economic profit last year= 150 000 + 5000×0.04= 150 000 + 200= $ 150 200